OpenStudy (anonymous):

Can someone please help me? I dont know how to work this problem! Perform the operation and leave the result in trigonometric form.

OpenStudy (anonymous):

\[\frac{ 4(\cos 40 \deg + isin40 \deg) }{ 3(\cos 160 \deg + isin 160 \deg) }\]